OpenStudy (anonymous):

Solve the equation. Check both solutions and only write the real solution(s). √2x+√3=x?

OpenStudy (anonymous):

Okay, square both sides. x^2 = 3 + 2x x^2 -2x-3 = 0 (x-3)(x+1) =0 So x = {3,-1} Now check. √2(3)+√3=3 √6+√3=3 Nope, so it's not 3. √2(-1)+√3=-1 √-2+√3=-1 √-2 is IMPOSSIBRU So no solutions?

OpenStudy (anonymous):

??? !! that's what i got to but on my paper there's like different choices

OpenStudy (anonymous):

I know, there's still a solution.

OpenStudy (anonymous):

A:(-1) B:(3) C:(-1,3) those are my choices ...

OpenStudy (anonymous):

Hm, maybe they don't want you to check? Then mark {3,-1}. That's what I got at first.

OpenStudy (anonymous):

ok then thank u ;P

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

it was wrong lol but thanks

OpenStudy (anonymous):

haha :P